16. Calculate the molarity of K+ ions in 0.76 g K3PO4 in 125 mL of solution
Answer #1


Molar mass of K3PO4,
MM = 3*MM(K) + 1*MM(P) + 4*MM(O)
= 3*39.1 + 1*30.97 + 4*16.0
= 212.27 g/mol


mass(K3PO4)= 0.76 g

use:
number of mol of K3PO4,
n = mass of K3PO4/molar mass of K3PO4
=(0.76 g)/(2.123*10^2 g/mol)
= 3.58*10^-3 mol
volume , V = 1.25*10^2 mL
= 0.125 L


use:
Molarity,
M = number of mol / volume in L
= 3.58*10^-3/0.125
= 2.864*10^-2 M

This is concentration of K3PO4
[K+] = 3*[K3PO4]
= 3*2.864*10^-2 M
= 8.59*10^-2 M
Answer: 8.59*10^-2 M